OpenStudy (anonymous):

what is the precision for a meter stick, vernier cailper, and micrometer caliper . what the relative uncertainity of an object of length 2 cm what is the the relative uncertainity is we use the above instruments

OpenStudy (anonymous):

meter stick = o.1cm vernier = 0.01 cm micrometer = 0.001cm relative uncertainity for 2cm will be for meter stick = (0.1/2 )*100% = 5% vernier = (0.01/2)*100% = 0.5% micrometer = (0.001/ 2)*100% = 0.05%